OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)?

OpenStudy (anonymous):

5-2/1-0?

OpenStudy (cwrw238):

- that the slope = 3

OpenStudy (anonymous):

So now i use y- y1 = m(x - x1)?

OpenStudy (cwrw238):

now you can use the general form y-y1 = m(x-x1) where m = 3 and (x1,y1) is one point on the line - say (0, 2)

OpenStudy (anonymous):

y- 2 = 3(x - 0)?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

I got that as an answer originally, but that wasn't even a choice on the assignment

OpenStudy (cwrw238):

in slope / intercept form this is y = 3x + 2

OpenStudy (anonymous):

the choices were y – 5 = 3(x – 1) y – 2 = –3(x – 0) y + 2 = 3(x + 1) y + 5 = –3(x + 1)

OpenStudy (anonymous):

y – 5 = 3(x – 1)

OpenStudy (precal):